OpenStudy (unimatix):

Prove that there is no positive rational number a such that a^2 = 3. You may assume that a positive integer can be written in one of the forms 3k, 3k + 1, 3k + 2 for some integer k. Prove that if the square of a positive integer is divisible by 3, then so is the integer. Then use a similar proof as for sqrt(2).

OpenStudy (faiqraees):

take square root on both side a = square root 3 a = 1.73 which is not a rational number

zepdrix (zepdrix):

Hmm I remember the sqrt2 proof :) That's a fun one.

OpenStudy (unimatix):

This is from Lang's Basic Mathematics by the way. So I'm looking for something around that level. Nothing too over my head.

ganeshie8 (ganeshie8):

This proof relies on fermats infinite descent

ganeshie8 (ganeshie8):

\(a^2 = 3\) Step1 : By way of contradiction, suppose that \(a\) is rational and is equal to \(\dfrac{x}{y}\) for some two "positive" integers. \(\left(\dfrac{x}{y}\right)^2 = 3\)

OpenStudy (unimatix):

Okay good! This looks similar to the stuff presented in the book.

ganeshie8 (ganeshie8):

Step2 : \(x^2=3y^2\) This means \(x^2\) is divisible by \(3\). It follows that \(x\) is also divisible by \(3\). Say, \(x = 3x_1\). Notice that \(\color{red}{x_1\text{ is less than }x}\). \((3x_1)^2 = 3y^2\)

ganeshie8 (ganeshie8):

good so far ?

OpenStudy (unimatix):

Yes. I was able to get steps 1 and 2 on my own, though I wasn't sure I was doing them correctly. After that is where it gets messy for me.

ganeshie8 (ganeshie8):

The logic in next steps is really beautiful

ganeshie8 (ganeshie8):

Again, this is proof by fermat's infinite descent. Has nothing to do with your textbook proof by contradiction

OpenStudy (unimatix):

Well, I don't know what that means, but hopefully I'll be able to follow it.

ganeshie8 (ganeshie8):

Step3 : \((3x_1)^2 = 3y^2 \\ \implies 3{x_1}^2=y^2\) This means \(y^2\) is divisible by \(3\). It follows that \(y\) is also divisible by \(3\). Say, \(y = 3y_1\). Notice that \(\color{red}{y_1\text{ is less than }y}\). \(3{x_1}^2 = (3y_1)^2\)

ganeshie8 (ganeshie8):

I'm sure you will get it easily :)

ganeshie8 (ganeshie8):

see if the first 3 steps make sense

OpenStudy (unimatix):

That makes sense!

ganeshie8 (ganeshie8):

What do you observe from first 3 steps

ganeshie8 (ganeshie8):

\(x_1\lt x\) and \(y_1\lt y\) yes ?

OpenStudy (unimatix):

Yes.

OpenStudy (unimatix):

That is true.

ganeshie8 (ganeshie8):

Good. Lets do two more steps before concluding

ganeshie8 (ganeshie8):

Step4 : \(3{x_1}^2 = (3y_1)^2 \implies {x_1}^2 = 3y_1^2\) This means \({x_1}^2\) is divisible by \(3\). It follows that \(x_1\) is also divisible by \(3\). Say, \(x_1 = 3x_2\). Notice that \(\color{red}{x_2\text{ is less than }x_1}\). \((3{x_2})^2 = 3{y_1}^2\)

ganeshie8 (ganeshie8):

Notice that \(x_2\lt x_1\lt x\)

ganeshie8 (ganeshie8):

Continuing the above steps, we get an infinite "decreasing" sequence of \(x_i\)'s and \(y_i\)'s

ganeshie8 (ganeshie8):

\[\ldots x_4\lt x_3\lt x_2\lt x_1\lt x\] and \[\ldots y_4\lt y_3\lt y_2\lt y_1\lt y\]

ganeshie8 (ganeshie8):

We're almost done. Still with me ? :)

OpenStudy (unimatix):

Sort of haha.

OpenStudy (unimatix):

It seems to make sense. Just kind of a weird.

OpenStudy (unimatix):

thing to do.

ganeshie8 (ganeshie8):

We're done actually. We just need to conclude

ganeshie8 (ganeshie8):

Before concluding, let me ask you a simple question.

ganeshie8 (ganeshie8):

what is your favorite large positive integer ?

OpenStudy (unimatix):

Sorry, I was going back and re-reading through those steps.

OpenStudy (unimatix):

5

ganeshie8 (ganeshie8):

that works, but pick something that is really large..

OpenStudy (unimatix):

55555

OpenStudy (unimatix):

Okay I'm confused on step 3 actually.

ganeshie8 (ganeshie8):

At the start of step3, I have just cancelled 3 both sides of the equation

OpenStudy (unimatix):

How did you do that on the left side?

OpenStudy (unimatix):

Never mind. I figured it out haha.

ganeshie8 (ganeshie8):

Nice :)

ganeshie8 (ganeshie8):

Now take your favorite integer : 55555

OpenStudy (unimatix):

So you would continue doing that same pattern forever, but how would that prove it?

ganeshie8 (ganeshie8):

Start with 55555, can you construct a "decreasing" sequence of positive integers ?

OpenStudy (unimatix):

Or rather how would that prove it's true.

ganeshie8 (ganeshie8):

I'm explaining you the exact same thing... :)

ganeshie8 (ganeshie8):

For example : ...., 3 < 100 < 4956 < 55555

ganeshie8 (ganeshie8):

Is that sequence really infinite ?

OpenStudy (unimatix):

Well no. Because eventually you would hit zero.

ganeshie8 (ganeshie8):

BINGO!

ganeshie8 (ganeshie8):

Starting with some finite positive integer, you can NEVER construct an "infinite" decreasing sequence of positive integers.

ganeshie8 (ganeshie8):

However we have got two "infinite" decreasing sequence of positive integers earlier : \[\ldots x_4\lt x_3\lt x_2\lt x_1\lt x\] and \[\ldots y_4\lt y_3\lt y_2\lt y_1\lt y\]

ganeshie8 (ganeshie8):

Since \(x\) and \(y\) are finite positive integers, above sequences cannot be infinite, yes ?

OpenStudy (unimatix):

WOAH.

ganeshie8 (ganeshie8):

Thats it! You have learned fermat's infinite descent !

ganeshie8 (ganeshie8):

We have arrived at a contradiction. Therefore we conclude that our initial assumption, \(a\) is rational, is wrong.

ganeshie8 (ganeshie8):

That ends the proof

OpenStudy (unimatix):

So that could be used on any similar problem involving a square right?

ganeshie8 (ganeshie8):

Absolutely! fermat's infinte descent hinges on this fact : The set of positive integers has a lower bound : 1. We cannot construct an "infinite" sequence of decreasing positive integers.

OpenStudy (unimatix):

But hold on, if you had say a^2 = 9 \[x^2 = 9y^2\] \[ x = 9x_{1}\] \[(9x_{1})^2 = 9y^2\] Then you could continue on for infinity right? And we know the sqrt(9) is 3.

ganeshie8 (ganeshie8):

You made a mistake. \(x^2\) is divisible by \(9\) need not imply that \(x\) is divisible by \(9\).

ganeshie8 (ganeshie8):

really good question though!

ganeshie8 (ganeshie8):

For example : \(3^2\) is divisible by \(9\). But \(3\) is not divisible by \(9\).

ganeshie8 (ganeshie8):

Your statement is true for primes though: If \(x^2\) is divisible by \(a\) and \(a\) is prime, then \(x\) is divisble by \(a\).

ganeshie8 (ganeshie8):

So you don't get an infinite sequence : \(x^2 = 9y^2\) we can only conclude that \(x=3x_1\) from above equation.

OpenStudy (unimatix):

But why is that?

ganeshie8 (ganeshie8):

what do you mean ?

OpenStudy (unimatix):

Ohhhh.

OpenStudy (unimatix):

I see the error.

ganeshie8 (ganeshie8):

Feel like doing another proof using fermat's infinite descent ?

OpenStudy (unimatix):

I'd love to go through another one.

ganeshie8 (ganeshie8):

let me find a neat problem..

ganeshie8 (ganeshie8):

Here it is : prove \(\large \sqrt[3]{2}\) is irrational

OpenStudy (unimatix):

I have another question though. The method given in the book to find prove a^2 = 2 involved showing that a/b was an even number over an even number when you tried to put it in the lowest form.

OpenStudy (unimatix):

And I was just wondering if you could use that kind of proof for an odd number. Because I was getting lost trying. I really like the Fermat's Infinite Descent though.

ganeshie8 (ganeshie8):

http://www.math.utah.edu/~pa/math/q1.html

ganeshie8 (ganeshie8):

That link has the proof that your textbook is talking about

OpenStudy (unimatix):

Yes. So that won't work for anything except for an even number?

ganeshie8 (ganeshie8):

It works for every irrational number

OpenStudy (unimatix):

Including the one we did earlier?

ganeshie8 (ganeshie8):

Yes, lets do a quick proof using simple contradiction for \(\sqrt{3}\)

OpenStudy (unimatix):

So \[a^3 = 2\] \[(\frac{ m }{ n })^3 = 2\]

OpenStudy (unimatix):

\[m^3 = 2n^3\]

ganeshie8 (ganeshie8):

are you trying fermat's infinite descent ?

OpenStudy (unimatix):

Oh I was going to. I guess you could use either?

OpenStudy (unimatix):

Sorry if that's a dumb question. I'm kind of new to anything involving proofs.

ganeshie8 (ganeshie8):

You could use either. Infinite descent is more powerful. Ppl don't use infinite descent for simple irrationalit proofs like these...

ganeshie8 (ganeshie8):

I have used infinite descent here just for some practice haha

ganeshie8 (ganeshie8):

Are you facing any problem proving the irrationality of \(\sqrt{3}\) using ur textbook method ?

OpenStudy (unimatix):

Yes. Hahaha.

OpenStudy (unimatix):

Though I feel like your method looks a bit more impressive haha. So \[a^2 = 3\] \[a = \frac{ m }{ n }\] \[m^2 = 3n^2\]

OpenStudy (unimatix):

Thus, m^2 is odd?

ganeshie8 (ganeshie8):

Thus, m^2 is divisible by 3

ganeshie8 (ganeshie8):

That means m is also divisible by 3, let m = 3k

OpenStudy (unimatix):

okay

ganeshie8 (ganeshie8):

plug that in the equation and get : (3k)^2 = 3n^2 3k^2 = n^2 This means n is also divisible by 3

ganeshie8 (ganeshie8):

So we see that both m and n are divisible by 3

ganeshie8 (ganeshie8):

But haven't you assumed in the start that m and n have no common factors ?